OpenStudy (anonymous):

It took Nora 36 hours to drive from Houston to Los Angeles. If the distance covered is 1,600 miles, what was Nora’s average speed in km per hour? You have to convert miles to km right? How?

OpenStudy (anonymous):

1600/36 = ?

OpenStudy (anonymous):

1m = 1.6 km. So, multiply that number by 1.6

OpenStudy (anonymous):

Which number?

OpenStudy (anonymous):

1600/36 times 1.6

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

I remeber doing something like : 1600 miles * 1km -------- = 1000miles or something similar ( i just can't remeber very well :( ) ^ Is this right?

OpenStudy (anonymous):

71 km/hr.

OpenStudy (anonymous):

GTTTTT! please tell me if it's right or wrong

OpenStudy (ksaimouli):

yup it is right

OpenStudy (anonymous):

So it becomes 1.6 km?

OpenStudy (ksaimouli):

nope 71

OpenStudy (ksaimouli):

km/hr

OpenStudy (ksaimouli):

convert 1m=1.6km

OpenStudy (ksaimouli):

just cross multiply

OpenStudy (anonymous):

No wait...

OpenStudy (anonymous):

I mean by 1.6 km when you convert

OpenStudy (anonymous):

not the final answer

OpenStudy (anonymous):

or

OpenStudy (anonymous):

71 is when you convert 1600miles to km?

OpenStudy (anonymous):

\[\Large \frac{1600 miles}{36 hours} \times \frac{1.6 km}{1 miles}=\frac{2560km}{36hours}\]

OpenStudy (anonymous):

YUUPPPP! This one :D

OpenStudy (anonymous):

But can you just tell me what you did exactly? Please :)

OpenStudy (anonymous):

It's a conversion factor. we know that there are 1.6km in a mile so you multiply the 1600 miles by 1.6km per 1mile the miles cancel out leaving 1600 miles times 1.6km the time portion is still in hours distance/time = rate

OpenStudy (anonymous):

\[\Large 1.6km = 1mile\] divide both sides by 1 mile \[\Large \frac{1.6km}{1mile}\]

OpenStudy (anonymous):

Okay so you did this: 1600 miles *1km/1000 miles = 1.6 km d=r*t t=36 hrs d= 1.6 km r= ? r= d/t r= 1.6/36 r=2/45 ?? ok i think i'm lost :(

OpenStudy (anonymous):

It took Nora 36 hours to drive from Houston to Los Angeles. If the distance covered is 1,600 miles, what was Nora’s average speed in km per hour? Where are you getting the 1km/1000 miles? 1km does not equal 1000 miles 1km = 0.62 miles or 1mile = 1.6km

OpenStudy (anonymous):

Ohhh! Okay Thankss :)